the fastest propeller - driven aircraft is the Russian TU- 95/142, which can reach a maximum speed of 925km/h for this speed, calculate the planes resultant displacement if it travels due east for 1.50h, then turns N45 degree East and travels for 2.0h (d= 1.32 × 10 rise to power of six meter, Φ= N3.5 degree East)
1
Expert's answer
2019-07-11T09:28:18-0400

Using law of cosines, square of total displacement is

"(925*1.5)^2+(925*2)^2-\newline 
-2*(925*1.5)*(925*2)*cos(180-45)="

"(925*1.5)^2+(925*2)^2-\newline 
-2*(925*1.5)*(925*2)*(-\sqrt{\smash[b]{2}})/2="

"(925*1.5)^2+(925*2)^2+\newline 
+(925*1.5)*(925*2)*(\sqrt{\smash[b]{2}})="

"925^2*(1.5^2+2^2+1.5*2*\sqrt{\smash[b]{2}})\approx"

"925^2*10.49264"

Total displacement is


"\sqrt{\smash[b]{925^2*10.49264}}=925*\sqrt{\smash[b]{10.49264}}\approx"

"925*3.239=2996.075"

Answer: 2996.075 km.